a baker purchased 12 lb of wheat flour and 15 lb of rye flour for a total cost of 18.30. a second purchase, at the same prices, included 15 lb of wheat flour and 10 lb of rye flour. the cost of the second purchase was 16.75. find the cost per pound of the wheat flour and of the rye flour.
Answer by pedjajov(51)

w - amount of wheat purchased
r - amount of rye purchased

12w+15r=18.30 /*2
15w+10r=16.75 /*3

24w+30r=36.60, subtract this equation from the other one
45w+30r=50.25

21w=13.65
w=13.65/21
w=0.65 is a cost per pound for wheat flour

From the first equation we have
r=(18.30-12w)/15
r=(18.30-12*0.65)/15
r=0.70 is a cost per pound for rye flour